# Exact non-stationary solutions of the Euler equations in two and three dimensions

### v1

@@ 4.1.2. Curl eigenfields in a circle bundle
$\operatorname{curl}|_{\mathrm{span}\{S_{i}\}}=\left(\begin{matrix}2\epsilon&\delta^{2}&0\\ 1&0&0\\ 0&in&0\end{matrix}\right),$
and it is easy to compute that the eigenvalues of this matrix are 0 and $\alpha_{\pm}$, with corresponding eigenvectors $z_{\pm}$. ∎
The Hopf fibration is the most well-known nontrivial circle bundle, mapping $S^{3}$ to $S^{2}$. In toroidal coordinates $(\chi,\theta,\phi)$ we can describe $S^{3}$ as a subset of $\mathbb{R}^{4}$ via
$x_{1}=\cos{\chi}\cos{\theta},\qquad x_{2}=\cos{\chi}\sin{\theta},\qquad x_{3}=\sin{\chi}\cos{\phi},\qquad x_{4}=\sin{\chi}\sin{\phi},$
where $\chi\in(0,\pi)$ and $\theta,\phi\in(0,2\pi)$. The metric in these coordinates is given by
$ds^{2}=d\chi^{2}+\cos^{2}{\chi}\,d\theta^{2}+\sin^{2}{\chi}\,d\phi^{2},$
and the Hopf vector field is given by the unit vector field $X=\partial_{\theta}+\partial_{\phi}$, with $2\pi$-periodic orbits. The volume form is given by
$\mu=\sin{\chi}\cos{\chi}\,d\chi\wedge d\theta\wedge d\phi,$
which fixes the orientation. Note that $X$ is a Killing field since the metric components do not depend on $\theta$ or $\phi$.
An oriented orthonormal basis of vector fields is given by
(4.9) $E_{1}=\partial_{\chi},\qquad E_{2}=\tan{\chi}\,\partial_{\theta}-\cot{\chi}\,\partial_{\phi},\qquad E_{3}=\partial_{\theta}+\partial_{\phi}=X,$
with dual basis of orthonormal $1$-forms given by
$E_{1}^{\flat}=d\chi,\qquad E_{2}^{\flat}=\sin{\chi}\cos{\chi}(d\theta-d\phi),\qquad E_{3}^{\flat}=\cos^{2}{\chi}\,d\theta+\sin^{2}{\chi}\,d\phi,$
satisfying $E_{1}^{\flat}\wedge E_{2}^{\flat}\wedge E_{3}^{\flat}=\mu$. Thus we can compute that
$dE_{3}^{\flat}=-2\sin{\chi}\cos{\chi}(d\chi\wedge d\theta+d\phi\wedge d\chi)=-2E_{1}^{\flat}\wedge E_{2}^{\flat}=-2\star\!E_{3}^{\flat},$
so that $\operatorname{curl}{X}=-2X$.
To find complex curl eigenfields satisfying the conditions of Theorem 2.2, we use Lemma 4.4 to reduce it to the question of complex eigenfunctions of the Laplacian, together with the condition $X(f)=inf$ for $n\in\mathbb{Z}$. Since $\partial_{\theta}$ and $\partial_{\phi}$ are both Killing fields, the Laplacian eigenfunctions must also be eigenfunctions of these operators, and the general form of $f$ will be
(4.10) $f(\chi,\theta,\phi)=e^{ij\theta}e^{ik\phi}\cos^{\lvert j\rvert}{\chi}\sin^{\lvert k\rvert}{\chi}J_{d}^{(\lvert k\rvert,\lvert j\rvert)}(\cos{2\chi})\qquad j,k\in\mathbb{Z},d\in\mathbb{Z}_{\geq 0}.$
Here we clearly have $n=j+k$, while $J_{d}^{(q,p)}$ is the Jacobi polynomial of degree $d$ given by
$J_{d}^{(q,p)}(x)=\frac{1}{2^{d}}\sum_{m=0}^{d}\binom{q+d}{m}\binom{p+d}{d-m}(x+1)^{m}(x-1)^{d-m}.$
See [19] for details.
Thus by Lemma 4.4, the curl eigenfields are obtained as follows. Given $j,k\in\mathbb{Z}$, $d\in\mathbb{Z}_{\geq 0}$, for
$\ell=\lvert j\rvert+\lvert k\rvert+2d,\qquad\alpha_{+}=\ell,\qquad\alpha_{-}=-(\ell+2),$
we find $\lambda$ in Theorem 2.2 by writing $\operatorname{curl}{X},z]=i\lambda\alpha_{\pm}z$, which here yields $\lambda=-2n/\alpha_{\pm}$. Thus for $f$ given by (4.10), we compute
$z_{\pm}=\Big(\alpha_{\pm}^{2}fE_{3}+\alpha_{\pm}\big(E_{2}(f)E_{1}-E_{1}(f)E_{2}\big)+i(j+k)\big(E_{1}(f)E_{1}+E_{2}(f)E_{2}+E_{3}(f)E_{3}\Big),$
in terms of the orthonormal basis (4.9).
The simplest special case is when $d=0$ with $j,k\geq 0$, when the Jacobi polynomial reduces to $J_{0}^{(\lvert k\rvert,\lvert j\rvert)}(x)=1$. In this case we see that $\ell=n=j+k$, and
$E_{1}(f)$ $=(k\cos^{2}{\chi}-j\sin^{2}{\chi})\xi(\chi,\theta,\phi)$
$E_{2}(f)$ $=-i(k\cos^{2}{\chi}-j\sin^{2}{\chi})\xi(\chi,\theta,\phi)$
$E_{3}(f)$ $=i(j+k)\sin{\chi}\cos{\chi}\xi(\chi,\theta,\phi)$
$\xi(\chi,\theta,\phi)$ $:=e^{ij\theta}e^{ik\phi}\cos^{j-1}{\chi}\sin^{k-1}{\chi}.$
We see that $z_{+}$ is identically zero in this case since $\alpha_{+}=j+k$, but $z_{-}$ is nontrivial, given by a multiple of
$z_{-}=\xi(\chi,\theta,\phi)\Big((k\cos^{2}{\chi}-j\sin^{2}{\chi})(iE_{1}+E_{2})+2(j+k)\sin{\chi}\cos{\chi}E_{3}\Big).$
The corresponding real solution of the $3D$ Euler equation on $S^{3}$ is given, using $\lambda-n=-\frac{n^{2}}{n+2}$ and any multiple $R$ of $z_{-}$, by
$u(t)=\partial_{\theta}+\partial_{\phi}+R\cos^{j-1}{\chi}\sin^{k-1}{\chi}\bigg[-(k\cos^{2}{\chi}-j\sin^{2}{\chi})\sin{\big(j\theta+k\phi-\tfrac{(j+k)^{2}}{j+k+2}t\big)}\,\partial_{\chi}\\ +\cos{\big(j\theta+k\phi-\tfrac{(j+k)^{2}}{j+k+2}t\big)}\bigg(\Big(3(j+k)\sin{\chi}\cos{\chi}-j\tan{\chi}\Big)\,\partial_{\theta}+\Big(3(j+k)\sin{\chi}\cos{\chi}-k\cot{\chi}\Big)\,\partial_{\phi}\bigg)\bigg].$
The apparent singularities in $\tan{\chi}$ and $\cot{\chi}$ are an artifact of the coordinate system; this is an analytic vector field on $S^{3}$, as one could see by writing everything in terms of the $(w,x,y,z)$ coordinates on $\mathbb{R}^{4}$ in terms of polynomials; again see [19].
For example in the simplest case where $j=1$ and $k=0$, we obtain
$u(t)=\partial_{\theta}+\partial_{\phi}+R\Big(\sin{\chi}\sin{(\theta-\tfrac{t}{3})}\,\partial_{\chi}+\cos{(\theta-\tfrac{t}{3})}\big((3\cos{\chi}-\sec{\chi})\,\partial_{\theta}+3\cos{\chi}\,\partial_{\phi}\big)\Big),$
which is the restriction to $S^{3}$ of the vector field on $\mathbb{R}^{4}$ given by
(4.11) $\begin{split}U(t)&=X+R\cos{(\tfrac{t}{3})}V_{1}+R\sin{(\tfrac{t}{3})}V_{2},\\ X&=-x_{2}\,\partial_{x_{1}}+{x_{1}}\,\partial_{x_{2}}-{x_{4}}\,\partial_{x_{3}}+{x_{3}}\,\partial_{x_{4}}\\ V_{1}&=-2{x_{1}}{x_{2}}\,\partial_{x_{1}}+(2{x_{1}}^{2}-{x_{3}}^{2}-{x_{4}}^{2})\,\partial_{x_{2}}+({x_{2}}{x_{3}}-3{x_{1}}{x_{4}})\,\partial_{x_{3}}+(3{x_{1}}{x_{3}}+{x_{2}}{x_{4}})\,\partial_{x_{4}}\\ V_{2}&=({x_{3}}^{2}+{x_{4}}^{2}-2{x_{2}}^{2})\,\partial_{x_{1}}+2{x_{1}}{x_{2}}\,\partial_{x_{2}}-({x_{1}}{x_{3}}+3{x_{2}}{x_{4}})\,\partial_{x_{3}}+(3{x_{2}}{x_{3}}-{x_{1}}{x_{4}})\,\partial_{x_{4}}.\end{split}$
This is essentially the simplest time-dependent solution one can construct on $S^{3}$ by our method, but there is an infinite-dimensional family of such solutions corresponding to different choices of $j$, $k$, and $d$, all of which will have polynomial coefficients in $\mathbb{R}^{4}$ and the same trigonometric time-dependence. Compare the standard Rossby-Haurwitz waves in Example 3.4. As with the standard ones, we may also take arbitrary linear combinations of vectors $z$ to get exact solutions as long as they have the same $n$ and $\ell$; that is, we can vary $j$, $k$, and $d$ as long as $j+k$ and $\lvert j\rvert+\lvert k\rvert+2d$ remain constant.
@@ 4.2. Torus Bundles over an Interval
Throughout this subsection we relabel the complex eigenfield from Theorem 2.2 as $v$, rather than $z$, to avoid conflict with the coordinate $z$ appearing in the adapted coordinate system.
@@ 4.2.1. The form of the metric
In any neighborhood where $X$ and $\omega$ are linearly independent, we may choose coordinates $(\theta,z)$ such that $X=\partial_{\theta}$ and $\omega=2\,\partial_{z}$, the coefficient chosen to make the formulas simpler. By formula (4.3), the function $\lvert X\rvert_{g}^{2}$ is not constant in this neighborhood. Take any nonsingular level surface $\lvert X\rvert=\text{constant}$ and construct geodesics which are initially unit and perpendicular to this surface, in the direction of $X\times\omega$. This generates a coordinate $r$ (the parameter along each geodesic) and a unit vector field $R=\partial_{r}$ satisfying $\nabla_{R}R=0$; see Petersen [28], Section 5.6. We claim that $g(R,X)\equiv 0$ and $g(R,\omega)\equiv 0$; indeed, since
$\frac{\partial}{\partial r}g(R,X)=g(\nabla_{R}R,X)+g(R,\nabla_{R}X)=0$
since $X$ is a Killing field, the fact that $g(R,X)=0$ on the initial surface implies that it always vanishes, and similarly for $g(R,\omega)$. We conclude that $g(\partial_{r},\partial_{\theta})=g(\partial_{r},\partial_{z})=0$. Furthermore by Lemma 4.1 we must have $g(\partial_{\theta},\partial_{z})=c$ for some constant $c$.
In the coordinates $(r,\theta,z)$ we can therefore write the metric components (which depend only on $r$) as
$ds^{2}=dr^{2}+\left(\varphi(r)\,d\theta+\frac{c}{\varphi(r)}\,dz\right)^{2}+\psi(r)^{2}\,dz^{2},$
for some functions $\varphi(r)$ and $\psi(r)$. An orthonormal basis of $1$-forms is given by
$(\xi^{1},\xi^{2},\xi^{3})=\left(dr,\varphi(r)\,d\theta+\frac{c}{\varphi(r)}\,dz,\psi(r)\,dz\right),$
and its dual orthonormal frame is given by
$(E_{1},E_{2},E_{3})=\left(\partial_{r},\frac{1}{\varphi(r)}\,\partial_{\theta},\frac{1}{\psi(r)}\,\Big(\partial_{z}-\frac{c}{\varphi(r)^{2}}\,\partial_{\theta}\Big)\right).$
We therefore have
$X^{\flat}$ $=\partial_{\theta}^{\flat}=\varphi(r)\,\xi^{2}=\varphi(r)^{2}\,d\theta+c\,dz$
$\omega^{\flat}$ $=2\,\partial_{z}^{\flat}=2\left(\psi(r)E_{3}+\frac{c}{\varphi(r)}\,E_{2}\right)^{\flat}=2\left(\psi(r)\xi^{3}+\frac{c}{\varphi(r)}\,\xi^{2}\right)$
@@
$dX^{\flat}=2\varphi(r)\varphi^{\prime}(r)\,dr\wedge d\theta=2\varphi^{\prime}(r)\left(\xi^{1}\wedge\xi^{2}+\frac{c}{\varphi(r)\psi(r)}\,\xi^{3}\wedge\xi^{1}\right)$
@@
$\star dX^{\flat}=2\varphi^{\prime}(r)\left(\xi^{3}+\frac{c}{\varphi(r)\psi(r)}\,\xi^{2}\right).$
But by definition of curl, we have $\star dX^{\flat}=\omega^{\flat}$, which implies $\psi(r)=\varphi^{\prime}(r)$ and yields (4.12). Performing the same computations in reverse, starting with the metric (4.12), we see that regardless of $c$ and $\varphi$, we do indeed get $\operatorname{curl}{\partial_{\theta}}=2\,\partial_{z}$, while both $\partial_{\theta}$ and $\partial_{z}$ are Killing fields since the metric components do not depend on either coordinate. Hence there are no further constraints. ∎
@@
In the special case where $c=0$ and $\varphi(r)=s_{k}(r)$ as in (4.4), this metric becomes
$ds^{2}=dr^{2}+s_{k}(r)^{2}\,d\theta^{2}+c_{k}(r)^{2}\,dz^{2},$
which is the standard constant curvature metric in polar coordinates for $k\in\{1,0,-1\}$ on $S^{3}$, $\mathbb{R}^{3}$, or hyperbolic $3$-space. Here the Killing field is the usual axisymmetric rotation $\partial_{\theta}$, which vanishes along the curve $r=0$. In this case we have $\operatorname{curl}{\partial_{\theta}}=2\partial_{z}$ while $\operatorname{curl}{\partial_{z}}=2k\partial_{\theta}$. When $k=1$ on the $3$-sphere, we see that $\partial_{\theta}+\partial_{z}$ is exactly the Hopf field from Example 4.5.
@@
so that for $c\neq 0$ we have a contact form. The corresponding Reeb field is $\xi=c^{-1}\,\partial_{z}$, which has squared length $\varphi^{\prime}(r)^{2}/c^{2}$. In the special case $\varphi(r)=r$ and $c=1$, the Reeb field is Killing and unit length, so we have a Sasakian geometry. The scalar curvature of the metric (4.12) in general is given by
@@
from which we see that the parameter $c$ genuinely changes the geometry, even in the special case $\varphi(r)=r$. For $\varphi(r)=\sin{r}$ we get the answer $R=6-2c^{2}\csc^{4}{r}$, showing that in general we cannot expect to make this metric well-defined at a point where $\varphi(r)=0$; hence it will typically only make sense on a noncompact manifold or a manifold with boundary. We do not know if there is any closed compact $3$-manifold that admits such a geometry with $c\neq 0$.
@@ 4.2.2. Curl eigenfields in a torus bundle
We now study the geometry of the $c$-metric (4.12), and in particular compute the Rossby-Haurwitz waves in this geometry. For concreteness suppose that $\theta$ and $z$ are both $2\pi$-periodic, so that $X=\partial_{\theta}$ and $\omega_{0}=2\partial_{z}$ span a torus for each fixed $r$ with closed trajectories, and assume $M$ is a solid torus bounded by $a\leq r\leq b$ with both $\varphi(r)$ and $\varphi^{\prime}(r)$ positive everywhere (which avoids potential singularities where either vanishes).
The following construction is very similar to the construction of curl eigenfields in Appendix C of [26], on the flat solid torus (with $c=0$ and $\varphi(r)=r$). Note that we cannot use the technique of Lemma 4.4 here, since although $\operatorname{curl}{(\partial_{\theta})}=2\,\partial_{z}$ by our definitions, the curl of $\partial_{z}$ does not simplify (see (4.13)), and in particular we would in general not get a finite-dimensional subspace by repeatedly taking curls.
@@
Vector fields satisfying the conditions of Theorem 2.2 on the 33-manifold $a,b]\times\mathbb{T}^{2}$ with metric (4.12) are given for integers $m,n$ by
(4.14) $v(r,\theta,z)=e^{in\theta}e^{imz}\bigg(i\,\frac{nh(r)-\big(m-\frac{cn}{\varphi(r)^{2}}\big)g(r)}{\alpha\varphi(r)\varphi^{\prime}(r)}\,\partial_{r}+\frac{g(r)}{\varphi(r)^{2}}\,\partial_{\theta}+\frac{h(r)}{\varphi^{\prime}(r)^{2}}\Big(-\frac{c}{\varphi(r)^{2}}\,\partial_{\theta}+\partial_{z}\Big)\bigg)$
where $g$ and $h$ solve the system
(4.15) $\alpha\varphi(r)\varphi^{\prime}(r)g^{\prime}(r)$ $=n\left(m-\frac{cn}{\varphi(r)^{2}}\right)\,g(r)+\big(\alpha^{2}\varphi(r)^{2}-n^{2}\big)h(r)$
(4.16) $\alpha\varphi(r)\varphi^{\prime}(r)h^{\prime}(r)$ $=\left(2c\alpha\,\frac{\varphi^{\prime}(r)^{2}}{\varphi(r)^{2}}+\left(m-\frac{cn}{\varphi(r)^{2}}\right)^{2}-\alpha^{2}\varphi^{\prime}(r)^{2}\right)\,g(r)-n\left(m-\frac{cn}{\varphi(r)^{2}}\right)\,h(r),$

### v2

@@ 4.1.2. Curl eigenfields in a circle bundle
$\operatorname{curl}|_{\mathrm{span}\{E_{i}\}}=\left(\begin{matrix}2\epsilon&\delta^{2}&0\\ 1&0&0\\ 0&i\zeta&0\end{matrix}\right),$
whose eigenvalues are $\alpha_{\pm}$ and 0 with corresponding eigenvectors $\mathfrak{q}_{\pm}$ and $E_{3}$ respectively. ∎
Remark 4.6.
Note that the existence of $f$ satisfying (4.5) comes from the fact that $X$ is a Killing field and hence commutes with the Laplacian. Hence, we can find an $L^{2}$ basis of simultaneous eigenfunctions. Each one generates an exact solution of the three-dimensional Euler equations via Theorem 2.2.
Armed with this, we now consider the flat and negatively curved analogues of Example 4.4.
Example 4.7 (The Heisenberg Manifold).
The Heisenberg group is the space of upper-triangular $3\times 3$ matrices with $1$ on the diagonal, which is homeomorphic to $\mathbb{R}^{3}$. The set of such matrices with integer entries is a discrete subgroup, and the quotient is a compact $3$-manifold called the Heisenberg manifold. It is a circle bundle over the flat torus $\mathbb{T}^{2}$. Eigenfunctions of the Laplacian can be worked out as in Tolimieri [38] using Hermite functions, and from these one could construct curl eigenfields in order to obtain explicit solutions. The formulas involve infinite sums, however, and are not nearly as elementary as in the case of the $3$-sphere, so we will not discuss them further here, although the interested reader could calculate them.
Example 4.8 (Hyperbolic Circle Bundles).
Let $\Sigma$ be a compact orientable surface of genus $g\geq 2$. By uniformization, $\Sigma$ admits a metric of constant curvature 1-1, realized as a quotient of the hyperbolic plane by a discrete group of Möbius transformations. In their study of the fast dynamo problem, Arnold and Khesin [2, Chapter V.4.D] illustrated that the unit tangent bundle of such a surface is a nontrivial circle bundle of the type appearing in Theorem 4.2, and carries natural curl eigenfields arising from this geometry. The same geometry gives examples that fit here, although the eigenfunctions of the Laplacian are harder to compute explicitly. The Euler class $k$ from Theorem 4.2 for a unit tangent bundle happens to coincide with the Euler characteristic, so $k=\chi(\Sigma)=2-2g$. The area of $\Sigma$ is given by the Gauss-Bonnet theorem as $\mathrm{Area}(\Sigma)=-2\pi\chi(\Sigma)=4\pi(g-1)$. Thus the curl eigenvalue is $2\epsilon=-1$. We refer the interested reader to [2] for details of the geometric construction. Note lastly that this construction also works over $S^{2}$, giving the bundle with Euler class $k=2$, but it does not work on $\mathbb{T}^{2}$ since the unit circle bundle is $\mathbb{T}^{3}$, which is a trivial bundle.
@@ 4.2.1. The form of the metric
In any neighborhood where $X$ and $\omega$ are linearly independent, they commute by (A.8), so we may choose coordinates $(\theta,z)$ such that $X=\partial_{\theta}$ and $\omega=2\,\partial_{z}$, where the normalization of $\omega$ is chosen for convenience. By formula (A.6), the function $\lvert X\rvert_{g}^{2}$ is not constant in this neighborhood. Let $\Sigma$ be a regular level surface of $\lvert X\rvert_{g}$. Since $X$ and $\omega$ are linearly independent, the vector field $X\times\omega$ is nowhere vanishing and orthogonal to both. We define a coordinate $r$ by taking unit-speed geodesics orthogonal to $\Sigma$ in the direction of $X\times\omega$. By construction,
$g(\partial_{r},\partial_{r})=1,\qquad\nabla_{\partial_{r}}\partial_{r}=0,$
cf. Petersen [31, Section 5.6]. We claim that $g(\partial_{r},X)=g(\partial_{r},\omega)=0$. Indeed,
$\partial_{r}\big(g(\partial_{r},X)\big)=g(\nabla_{\partial_{r}}\partial_{r},X)+g(\partial_{r},\nabla_{\partial_{r}}X).$
The first term vanishes since $\nabla_{\partial_{r}}\partial_{r}=0$, and the second vanishes because $X$ is Killing. Since $g(\partial_{r},X)=0$ along $\Sigma$, it follows that $g(\partial_{r},X)\equiv 0$. The same argument applies to $\omega$. Hence
$g(\partial_{r},\partial_{\theta})=g(\partial_{r},\partial_{z})=0.$
By Proposition 4.1, the quantity $g(\partial_{\theta},\partial_{z})$ is constant; denote it by $c$. Since the metric is invariant under both Killing fields, its coefficients depend only on $r$. It follows that the metric can be written in the form
$ds^{2}=dr^{2}+\left(\varphi(r)d\theta+\frac{c}{\varphi(r)}dz\right)^{2}+\varsigma(r)^{2}dz^{2},$
for positive functions $\varphi(r)$ and $\varsigma(r)$. An orthonormal basis of $1$-forms is given by
$\xi^{1}=dr,\quad\xi^{2}=\varphi(r)\,d\theta+\frac{c}{\varphi(r)}\,dz,\quad\xi^{3}=\varsigma(r)\,dz.$
@@
$X^{\flat}=\partial_{\theta}^{\flat}=\varphi(r)^{2}\,d\theta+c\,dz=\varphi(r)\,\xi^{2},$
and hence
$dX^{\flat}=2\varphi(r)\varphi^{\prime}(r)\,dr\wedge d\theta.$
Expressing this in the orthonormal coframe gives
$dX^{\flat}=2\varphi^{\prime}(r)\left(\xi^{1}\wedge\xi^{2}+\frac{c}{\varphi(r)\varsigma(r)}\,\xi^{3}\wedge\xi^{1}\right),$
@@
$\star dX^{\flat}=2\varphi^{\prime}(r)\left(\xi^{3}+\frac{c}{\varphi(r)\varsigma(r)}\,\xi^{2}\right).$
On the other hand,
$\omega^{\flat}=2\,\partial_{z}^{\flat}=2\left(\varsigma(r)\,\xi^{3}+\frac{c}{\varphi(r)}\,\xi^{2}\right).$
Comparing $\star dX^{\flat}=\omega^{\flat}$ yields $\varsigma(r)=\varphi^{\prime}(r)$, and substituting this into the metric gives (4.7).
Conversely, for any constant $c$ and positive function $\varphi$, the vector field $\partial_{\theta}$ is Killing, as the metric components are independent of $\theta$. We can then compute that $\operatorname{curl}(\partial_{\theta})=2\,\partial_{z}$ which, since the metric coefficients are also independent of $z$, is a linearly independent Killing field. This completes the proof. ∎
@@
In the special case where $c=0$ and $\varphi(r)=s_{\kappa}(r)$ as in (3.1), this metric becomes
$ds^{2}=dr^{2}+s_{\kappa}(r)^{2}\,d\theta^{2}+s_{\kappa}^{\prime}(r)^{2}\,dz^{2},$
which is the standard constant curvature metric in polar coordinates for $\kappa\in\{1,0,-1\}$ on $S^{3}$, $\mathbb{R}^{3}$, or hyperbolic $3$-space. Here the Killing field is the usual axisymmetric rotation $\partial_{\theta}$, which vanishes along the curve $r=0$. In this case we have $\operatorname{curl}{\partial_{\theta}}=2\partial_{z}$ while $\operatorname{curl}{\partial_{z}}=2\kappa\partial_{\theta}$. When $\kappa=1$ on the $3$-sphere, we see that $\partial_{\theta}+\partial_{z}$ is exactly the Hopf field from Example 4.4. Hence the formulas here may be used to obtain new solutions analogous to the ones presented there.
@@
so that for $c\neq 0$ we have a contact form. The corresponding Reeb field is $\xi=c^{-1}\,\partial_{z}$, which has squared length $\frac{\varphi^{\prime}(r)^{2}}{c^{2}}$. In the special case $\varphi(r)=r$ and $c=1$, the Reeb field is Killing and unit length, so we have a Sasakian geometry. The scalar curvature of the metric (4.7) in general is given by
@@
from which we see that the parameter $c$ genuinely changes the geometry, even in the special case $\varphi(r)=r$. For $\varphi(r)=\sin{r}$ we get $R=6-2c^{2}\csc^{4}{r}$, showing that in general we cannot expect to make this metric well-defined at a point where $\varphi(r)=0$; hence it will typically only make sense on a non-compact manifold or a manifold with boundary. We do not know if there is any closed $3$-manifold that admits such a geometry with $c\neq 0$.
@@ 4.2.2. Curl eigenfields in a torus bundle
We now study the geometry of the $c$-metric (4.7), and construct solutions to the Euler equations (1.1) in this geometry via Theorem 2.2. For concreteness suppose that $\theta$ and $z$ are both $2\pi$-periodic, so that $X=\partial_{\theta}$ and $\omega_{0}=2\partial_{z}$ span a torus for each fixed $r$ with closed trajectories, and assume $M$ is a solid torus bounded by ${\varepsilon_{1}}\leq r\leq{\varepsilon_{2}}$ with both $\varphi(r)$ and $\varphi^{\prime}(r)$ positive everywhere (which avoids potential singularities where either vanishes).
The following construction is similar to that of the curl eigenfields in Appendix C of [29], on the flat solid torus (with $c=0$ and $\varphi(r)=r$). Note that we cannot use the technique of Lemma 4.5 here, since although $\operatorname{curl}{(\partial_{\theta})}=2\,\partial_{z}$ by our definitions, the curl of $\partial_{z}$ does not simplify, cf. (4.8). In particular we would in general not get a finite-dimensional subspace by repeatedly taking curls. Therefore the computations need to be done directly as in the proof of Theorem 4.9. The details of the proof are presented in Appendix A.
@@
Consider the three-dimensional manifold ${\varepsilon_{1}},{\varepsilon_{2}}]\times\mathbb{T}^{2}$ with metric (4.7). For $m,n\in\mathbb{Z}$, let $a$ and $b$ solve the Sturm-Liouville type system
(4.9) $\alpha\varphi(r)\varphi^{\prime}(r)a^{\prime}(r)$ $=n\left(m-\frac{cn}{\varphi(r)^{2}}\right)\,a(r)+\big(\alpha^{2}\varphi(r)^{2}-n^{2}\big)b(r)$
(4.10) $\alpha\varphi(r)\varphi^{\prime}(r)b^{\prime}(r)$ $=\left(2c\alpha\,\frac{\varphi^{\prime}(r)^{2}}{\varphi(r)^{2}}+\left(m-\frac{cn}{\varphi(r)^{2}}\right)^{2}-\alpha^{2}\varphi^{\prime}(r)^{2}\right)\,a(r)-n\left(m-\frac{cn}{\varphi(r)^{2}}\right)\,b(r),$